Solve using the five-step problem-solving process. Show all steps necessary to arrive at your solution.
A cement walk of uniform width is built around a 20 foot by 40 foot rectangular pool and the area of the walk is 700 square feet. How wide will the walk be?

Answer by richwmiller(17219)

the pool has an area of
20*40=800
(20+2x)*(40+2x)=1500
positive answer only
x=5
the walkway will be 5 feet wide
check
30*50=1500
ok
why 2x? because the path is on both sides of the pool